Two snails, Alfie and Brian, are 1 meter apart and start sliding towards each other at the same time. Alfie slides 8mm every minute. Brian slides 12mm every minute.
a) After how many minutes will the snails meet?
b) How much further than Alfie will Brian have moved?

Answer by ankor@dixie-net.com(22740)

Two snails, Alfie and Brian, are 1 meter apart and start sliding towards each other at the same time.
Alfie slides 8mm every minute. Brian slides 12mm every minute.
:
Change 1 meter to 1000 mm
let t = travel time in minutes when they meet
a) After how many minutes will the snails meet?
write a distance equation; dist = time * speed
8t + 12t = 1000
20t = 1000
t = 1000/20
t = 50 min
b) How much further than Alfie will Brian have moved?
50 * 12 = 600 mm
50 * 8 = 400 mm
------------------
dist dif: 200 mm